# HCT-QA: A Benchmark for Question Answering on Human-Centric Tables

### v2

@@ Abstract
Tabular data embedded in PDF files, web pages, and other document types is prevalent in different sectors such as government, science, and business. These tables, which we call human-centric tables (HCTs), are dense in information but often exhibit both structurally and semantically complex layouts. This complexity poses significant challenges to querying them using traditional data processing systems. While some existing solutions focus on transforming these tables into relational formats, they fail to handle the diverse and complex layouts of HCTs, making them not amenable to easy querying with SQL-based approaches. Another promising option is to use Large Language Models (LLMs) and Vision Language Models (VLMs). To settle this debate and make progress on answering queries on this type of table, we propose HCT-QA, an extensive benchmark of HCTs, natural language queries, and their respective answers on thousands of tables. Our dataset includes 1,880 real-world HCTs with 9,835 QA pairs in addition to 4,679 synthetic HCTs with 67.7K QA pairs. Furthermore, we show through a series of extensive experiments the performance of a large number of LLMs (21) and VLMs (13) in properly answering the different questions in our benchmark. We also show how fine-tuned models can significantly outperform their pretrained versions on this task.
1 Introduction
Tabular data embedded in PDF files, web pages, and other types of documents are commonly present in numerous domains. Figure 1 shows a few examples of such data obtained from the US Census (uscensus, ), the Pakistan Bureau of Statistics (pakcensus, ), the Qatar National Planning Council (qnpc, ), and scientific paper archives (arxiv, ). These tabular data representations have rich semantics and are usually used to make strategic decisions without revisiting the raw, low-level data, often unavailable to the end user. They represent a type of data modality that encompasses a unique set of characteristics that make them very valuable to businesses, but extremely challenging to process systematically.
Challenging layouts.
The data is visualized in a non-standard, often complex tabular layout with artistic adornment to deliver a wealth of information summarized in a concise and dense way targeted primarily for direct human understanding and consumption. In Figure 1, we label several of the common table layout properties for ease of reference throughout the paper, including column nesting (L1), row nesting (L2) or row group labels (L3), embedded aggregations either row-wise (L4) or column-wise (L5). These properties may rely on visual traits such as differentiating text or row colors (L3 in table (d) & L4 in table (c)), bold text (L4 in tables (a) & (c)), or positioning within the cell to carry some semantic meaning (L2 in table (b)). Some other layout properties include nested table cells, local aggregations for specific subset of columns or rows, cell dividers, and row-group clustering (See examples in Appendix A). Although suitable for humans, such complex layouts introduce numerous challenges for systematic and automated methods for processing and querying the data. For instance, these layouts make it challenging to automatically answer questions like “What is the amount of dates imported in 2015?” (Table (a)), or “Which subject had the highest number of hardcover books sold in 2000” (Table (c)).
Untapped huge potential.
Tabular data embedded in documents are often treated as terminal data, i.e., no further processing is expected beyond eyeball analysis, which limits their full potential, especially at scale. For example, a single organization might have tens of thousands of such tables, and a big data lake could hold millions eltabakh2023crossmodaldatadiscovery . Moreover, for many applications, these tabular representations (and their container documents) are the only source of such aggregated data. Indeed, the base data (e.g., relational databases, CSVs, or Excel sheets) may not be available due to undocumented or lost sources, restricted access to the underlying databases, or missing ETL generating processes. Additionally, the content in these tables could be computed and assembled from multiple disjoint sources and augmented with external domain-expert knowledge. Hence, such data is unique and highly valuable but hardly exploitable at scale. In the rest of the paper, we refer to this tabular modality intended primarily for human consumption as “Human Centric Tables” ((HCTs for short)).
@@ 1 Introduction
Figure 1: Examples of Human-Centric Tables (HCTs) from different real-world data sources exhibiting various levels of structural and semantic complexity: column nesting (L1); row nesting (L2); row group labels (L3); row aggregation (L4); column aggregation (L5).
Although tools exist to extract HCTs from PDF documents or web pages aspose ; innodata ; pdftable-tableextractor ; pdfplumber , end-users are on their own with no systematic support for any further processing or analytics. Research works typically study ways to transform non-relational layouts into relational tables using heuristics or data-driven transformation techniques AutoTables_57 ; retclean-paper ; tabnet-paper ; tab-cleaner-paper , then use NL-to-SQL approaches NLtoSQLsurvey2024 to query them. This process is complex, costly, error-prone, and may not work for elaborate HCTs.
A new perspective.
@@
Recent advances in Large Language Models (LLMs) and Vision Language Models (VLMs) motivate a new “LLM-oriented” approach for querying HCTs, which could (1) efficiently handle HCTs among other human-centric data formats (images, charts, videos…), (2) avoid the complex and error-prone NL-to-SQL pipeline, and (3) directly support natural language queries, particularly crucial for multilingual tables such as the ones from the Qatar National Planning Council (Figure 1a). However, a systematic evaluation for identifying the strengths, weaknesses, and potential improvements of LLM-oriented approaches remains uncharted. Hence, we propose a novel benchmark for assessing LLMs as the engine for understanding, processing, and querying HCTs.
@@
Challenges and contributions. Devising the proposed benchmark involves several challenges: (1) collecting or generating HCTs with an extensive diversity of structures from relational table layouts to very complex layouts as illustrated in Figure 1; (2) collecting or generating an extensive variety of natural language questions that require steps from simple selection to complex aggregation, joins, and ranking questions; (3) getting complete and accurate ground truth answers to each table and question pair; and (4) exploring strategies like fine-tuning and impact on performance. Here are our key contributions:
@@
  • We propose a large HCT and question-answer (QA) benchmark containing 1,880 HCTs from four real-world data sources spanning over fifty semantic domains, both as images and extracted CSV formats, along with 9,835 QA pairs crafted by experts.
@@
  • We develop a novel synthetic HCT generator to complement the real-world datasets with 4,679 HCTs, along with 67,747 template-based QA pairs covering seven semantic domains.
  • We propose layout and question properties capturing the most common elements of HCTs and natural language questions to identify performance gaps and opportunities for improvement of processing engines.
@@
  • We use our benchmark and properties to evaluate $21$ LLMs and $13$ VLMs ranging from $3B$ to $100B+$ parameters. Our evaluation reveals that large closed-weight models perform the best overall, gpt-o1-mini being the best-performing model with a Recall score of 68.5%. Also, the small Pixtral-12B VLM outperforms the best Qwen2.5-7B LLM of comparable size by 7 points in Recall, showing promise of processing HCT images directly.
  • We fine-tuned two models and showed that fine-tuning is indeed effective and can improve performance by up to 50 points compared to the pretrained versions.
@@
Our benchmark is publicly available on HuggingFace (HCT_QA_hf, ). All experimentation codes and the synthetic HCT generator are publicly available on our GitHub repository (HCT_QA_github, ) .
2 Related work
There is a flurry of research and commercial systems on document understanding and information extraction from complex documents (especially PDF) Ahmed2016_56 . However, none of this work tries to answer the questions targeted in the proposed benchmark, i.e., the potential and the hurdles that face the LLM-based approach as the query engine for HCTs.
Table extraction techniques transform tables from PDF documents or images into machine-readable format like CSV, HTML, and XML. There are various commercial tools  (aspose, ; azure_document_intelligence, ; innodata, ; docugami, ) and recent literature (tableformer_2022, ; robustTableDetection_Ma_2022, ; omniParser_2024, ; smolDocling_2025, ; Khan2021ICDAR_13, ; Oro2008ICDAR_14, ; Qiao2021ICDAR_22, ) that show promise on extracting tables like HCTs into these formats. Nevertheless, beyond the extraction step, end-users are on their own with no systematic support for further processing, querying, analytics, or discovery.
NL-to-SQL Weir2020ACM_25 ; Basik2018ACM_26 ; Wang2019_27 ; Brunner2021ValueNetAN_28 ; sen2020VLDB_29 ; liu2024surveynl2sqllargelanguage These techniques assume clean and simple relational tables as input and focus on translating natural language queries to SQL. Several techniques are proposed to transform non-relational tabular data into relational tables 10.1431751.3231766 ; patsql_vldb2021 ; 10.147407790.3407831 ; AutoTables_57 . However, these techniques deal with much simpler layouts compared to the HCTs presented in Figure 1, and the transformation accuracy reported in the most recent work AutoTables_57 is about 60%.
Existing Table QA. Most existing Table QA benchmarks use flat Wikipedia tables which, when compared to HCTs, are significantly simpler to understand tablebench_2025 ; wikitablequestions_2017 ; fetaqa_2021 ; tabfact_2020 ; ottqa_2021 ; hybridqa_2021 . Other benchmarks use tables from one or two semantic domains and use a very small number of tables.  aitqa_2021 ; TAT-QA ; finqa_2022 ; scitat_2024 ; multitat_2025 ; tablevqa_2024 ; spiqa_neurips2024 . Some benchmarks focus on QA on documents that include different modalities such as tables, charts, and text paragraphs but not on isolated HCTs TAT-QA ; scitat_2024 ; MMLongbenchdoc_neurips2024 . In comparison, HCT-QA uses 1,880 real HCTs from 4 data sources and 4,679 synthetic HCTs that cover 50+ domains, and have questions that require only the HCT as the context to be answered.
3 HCT-QA benchmark datasets
In this section we describe the construction of our benchmark datasets. Each element in the datasets consists of a triplet in the form of $(T_{HCT},Q_{NL},A)$ along with rich metadata for facilitating deeper analysis. We focus on the real-world HCT datasets collected from thousands of PDF documents from diverse domains. We present the extraction and transformation processes that we applied to them, and the most common table patterns present in such data. We then introduce our proposed characterization of query properties. Finally we present the synthetic table and QA generator. Further details are given in Appendix A,  Appendix B, and Appendix C
Overall, our benchmark consists of real-world and synthetic tables and QA summarized in Figure 2.
3.1 Real-world dataset collection
Data sources:
@@ 3.1 Real-world dataset collection
We collected tens of thousands of PDF documents from diverse sources: Qatar National Planning Council (QNPC) qnpc (Figure 1(a)), ArXiv Research Papers from ArXiv.org, biorxiv.org, and medarch.org (Figure 1(b)), US Census covering PDF reports from the US Census Bureau uscensus (Figure 1(c)) and Pakistan Census from the official website of the Pakistan government pakcensus (Figure 1(d)). These sources cover more than 50 domains ranging from education, disabilities, land usage, capital gain, machine learning, vision, data management, finance, astrophysics, nanoscale and nuclear physics, materials science, phenomenology, to cryptography and security, lipid disorders, hypertension, diabetology, preventive cardiology, and many others.
Extraction and transformation:
We propose a canonical representation of HCTs in CSV format wherein (i) for column nesting (L1) all parent cell values are above the left most child cell, (ii) for row nesting (L2) all parent cell values are in the same row as the top-most child cell, (iii) row group labels (L3) are added in the left most cell of that row with the remaining row cells as blank (L3 $\rightarrow$ L2), and (iv) single cell values in borderless tables that span multiple lines are represented only in the cell that corresponds to the top line (see Appendix E for illustrative examples). Other properties, such as aggregation rows (L4) and columns (L5) are explicitly represented as regular well-formatted rows and columns in the CSV. This representation in CSV format allows for easy understanding of the table hierarchy and enables further rule-based transformations on the HCT if needed.
After exploring various open-source and commercial tools such as Tabula (tabula-py, ), PDF Plumber pdfplumber , Azure AI Document Intelligence azure_document_intelligence , Aspose aspose , Docling smolDocling_2025 ; doclingModels_2025 , CascadeTabNet (cascadeTabNet_2020, ), we found that Azure AI Document Intelligence was satisfactory in its consistency at outputting the HCT image in PNG or JPG format and its canonical CSV representation, and being the cheapest and most convenient to use as an API service. See Appendix E for details on the extraction performance.
Characterization of table properties:
Across domains and sources, we found that certain structural layouts are very dominant (e.g., column nesting, and row nesting) compared to rare and outlier patterns (e.g., row group labels that are also local row aggregations). In the current benchmark, we focused only on the common patterns and excluded tables exhibiting outlier patterns. We determined properties to characterize HCTs related to the number of levels and symmetry of column and row headers (L1, L2, L3), the presence of aggregate rows or columns (L4, L5) explicitly labeled or not, and the indentation structure to display them, as well as split lines or color adornments to distinguish semantically similar areas in the table. Figure 2 gives the distribution of main Real-World HCT structural properties (a,d). The full list of properties and the web interface we developed to annotate them manually are described in Appendix A.
@@
(a) Statistics of HCTs and QA pairs (b) Main properties of real-world QA pairs

### v3

# Mohammad S. Ahmad    Zan Naeem*    Michael Aupetit    Ahmed Elmagarmid    Mohamed Eltabakh    Xiaosong Ma* * Work done while at QCRI.    Mourad Ouzzani    Chaoyi Ruan*    Hani Al-Sayeh

@@ Abstract
Tabular data embedded in PDF files, web pages, and other types of documents is prevalent in various domains. These tables, which we call human-centric tables (HCTs for short), are dense in information but often exhibit complex structural and semantic layouts. To query these HCTs, some existing solutions focus on transforming them into relational formats. However, they fail to handle the diverse and complex layouts of HCTs, making them not amenable to easy querying with SQL-based approaches. Another emerging option is to use Large Language Models (LLMs) and Vision Language Models (VLMs). However, there is a lack of standard evaluation benchmarks to measure and compare the performance of models to query HCTs using natural language. To address this gap, we propose the Human-Centric Tables Question-Answering extensive benchmark (HCT-QA) consisting of thousands of HCTs with several thousands of natural language questions with their respective answers. More specifically, HCT-QA includes 1,880 real-world HCTs with 9,835 QA pairs in addition to 4,679 synthetic HCTs with 67.7K QA pairs. Also, we show through extensive experiments the performance of 25 and 9 different LLMS and VLMs, respectively, in an answering HCT-QA’s questions. In addition, we show how finetuning an LLM on HCT-QA improves F1 scores by up to 25 percentage points compared to the off-the-shelf model. Compared to existing benchmarks, HCT-QA stands out for its broad complexity and diversity of covered HCTs and generated questions, its comprehensive metadata enabling deeper insight and analysis, and its novel synthetic data and QA generator.
I Introduction
Answering natural language (NL) questions over structured and semi-structured data is a fundamental problem [62, 56, 1, 49]. The task of Table QA focuses on retrieving or computing answers to NL questions grounded in tabular data [21, 65, 6, 16, 61, 17]. Early studies used primarily relational tables from Wikipedia and widely used databases [65, 33, 23]. Over time, research has evolved in two directions: (1) Direct question answering using neural models [21, 37, 25], and (2) translating NL questions into SQL queries (NL-to-SQL) [32, 6, 16, 61, 17]. Despite increasing attention to this area, existing efforts remain concentrated on relational or relational-like table structures.
Tabular data embedded in PDFs files, web pages, and other types of documents is widespread across many domains. Figure 1 illustrates examples of such data obtained from the Qatar National Planning Council [51], scientific publication archives [2], the US Census [58], and the Pakistan Bureau of Statistics [48]. These tables employ intricate structural and visual layouts to convey large volumes of information in a concise representation to support strategic decision-making.
@@ I Introduction
Figure 1: Examples of Human-Centric Tables (HCTs) from different real-world data sources exhibiting various levels of structural and semantic complexity including column nesting (L1), row nesting (L2), row group labels (L3), row aggregation (L4), and column aggregation (L5)
Unlike traditional relational tables, which have a simple layout, document-embedded tables are primarily designed for human readability rather than automated processing. We refer to this form of tabular data as Human-Centric Tables (HCTs). HCTs represent a distinct data modality, characterized by intricate layouts, rich semantic content, and significant practical value. However, systematically processing HCTs for tasks such as information extraction, interpretation, and question answering remains a major challenge.
With the rapid advances in multimodal document understanding [55, 36, 39, 35], the scope of Table QA is progressively extending beyond relational tables [17, 16, 42]. This evolution warrants the need for not only systems capable of robust question answering over HCTs, but also methodologies and benchmarks that enable rigorous evaluation and identification of existing limitations of these systems while answering questions on HCTs. Although a few benchmarks have been recently introduced, they suffer from key shortcomings—most notably, the absence of metadata-driven analyses (i.e., analyses incorporating HCTs and question characteristics), and limited scalability due to the reliance on extensive manual effort.
Properties and Challenges of HCTs. HCTs exhibit highly diverse structural features designed for human interpretation. Figure 1 showcases common structural properties observed in HCTs including column nesting (L1), row nesting (L2), and row group labels (L3), as well as embedded aggregations that may appear row-wise (L4) or column-wise (L5). Visual cues-such as variations in text or row color (e.g., L3 in Table (d) and L4 in Table (c)), bold or emphasized text (e.g., L4 in Table (a) and Table (c)), and the positioning of elements within cells (e.g., L2 in Table (b)) convey additional semantic information.
These structural and visual complexities make direct question answering with neural QA models challenging. Also, as HCTs are not inherently relational, SQL queries cannot be executed on them directly. Existing efforts, such as AutoTables [34], have attempted to transform HCT-like tables into relational formats. However, such approaches are limited to relatively simple layouts, often struggle on complex real-world HCTs. Consequently, the pipeline that first converts HCTs to relational tables and then applies NL to SQL transformations remains fragile and error-prone [29, 17, 16, 42].
@@
New Opportunities with LLMs and VLMs. Recent advances in Large Language Models (LLMs) and Vision Language Models (VLMs) pave the way for a new LLM-oriented approach for querying HCTs. This approach can (1) efficiently handle HCTs alongside other human-centric data formats (e.g., charts and images), (2) avoid the complex and error-prone NL-to-SQL pipeline, and (3) directly understand NL questions. However, a systematic evaluation to identify the strengths, weaknesses, and potential improvements of LLM-oriented approaches remains largely an uncharted area. To address this gap, we propose HCT-QA, a novel benchmark designed to assess LLMs and VLMs as the engines for querying HCTs. HCT-QA can be also used to evaluated other types of data processing engines designed to answer NL questions on HCTs.
TABLE I: Comparison of major Table QA benchmarks with HCT-QA.
Benchmark Name Data Type Num HCT Tables Table Modalities Num Sources Num QA Pairs Table Metadata QA Metadata Avg. Questions Per Table Question Syntactic Similarity
HiTab [13] Tables 3,597 JSON 3 10,672 3 0.45
MultiHiertt [69] Tables + Text 9,842 HTML 2 10,440 N/A 0.23
TableVQA-Bench [30] Tables 205 HTML, Image 1 250 1.2 0.186
TAT-QA [71] Tables + Text 2,757 2D List 1 12,650 N/A 0.2879
HCT-QA (Ours) Tables 6,559 HTML, CSV, Image, MD 4 + Synthetic 77,582 11 0.325
@@
Challenges and Contributions. Devising the proposed benchmark involves several challenges: (1) Collecting and generating HCTs with an extensive diversity of complex layouts (Figure 1). (2) Constructing an extensive variety of NL questions ranging from simple selection to complex aggregation and ranking questions. (3) Obtaining accurate ground truth answers, and developing meaningful evaluation metrics to compare them to LLMs’ answers. (4) Exploring the impact of LLM-specific strategies such as finetuning on answering questions on HCTs. The key contributions of HCT-QA are:
@@
  • A large set of HCTs and question-answer pairs (QA) consisting of 1,880 HCTs from four real-world data sources spanning diverse domains. HCTs are provided in various formats including images, CSV, HTML, and Markdown. We also provide 9,835 QA pairs crafted by experts and verified by annotators (Section III-A).
  • Extensive manually-obtained metadata of the HCTs and QA pairs to capture the most common properties of HCTs and questions. This metadata allows for a deeper understanding of LLM/VLM performance and a better identification of performance gaps (Section III-B).
@@
  • A novel synthetic HCT generator to complement the real-world datasets. We used the generator to produce 4,679 HCTs and 67,747 template-based QA pairs across seven semantic domains. The generator offers flexible configuration making it a valuable and practical resource for the research community (Section IV).
@@
  • An extensive evaluation of $25$ LLMs and $9$ VLMs of different sizes ranging from $3B$ to $100B+$ parameters. Our evaluation reveals that large closed-weight models perform the best overall, the promise of processing HCT images directly using VLMs, and the effectiveness of finetuning some models using HCT-QA (Section V).
@@
Our benchmark is publicly available on HuggingFace [19]. All experimentation code and the synthetic HCT generator are publicly available on our GitHub repository [18].
II Related work
Several research and commercial efforts addressed document understanding and information extraction from complex documents [55, 36, 35]. However, none of these fully address the core question What are the opportunities and challenges of using LLM-based approaches to query HCTs?
Table extraction techniques in several commercial tools [3, 14, 22, 15] and recent research [7, 44, 40, 59, 45, 28, 47, 52] focus on extracting tables similar to HCTs from PDFs into machine-readable formats such as CSV and Markdown. However, they do not provide end-users any systematic support for querying, analyzing, or discovering the extracted tables. Also, our analysis shows that VLMs can bypass this extraction step by operating directly on the image version of HCTs.
Table QA approaches focus on relational tables as they use NL-to-SQL to answer the questions [63, 4, 60, 5, 54, 38]. However, SQL queries cannot be executed on tables with complex layouts like HCTs. As mentioned above, AutoTables [34] fail to achieve high accuracy on complex real-world HCTs. Other approaches tried to perform repairs and transformations on non-relational tables but still do not present a means to convert a HCT into a relational table [20, 26]. Such a multi-step, database-oriented pipeline-spanning HCTs extraction, schema discovery, and NL-to-SQL transformation introduces several error-prone phases and could be an overkill for practical applications. Other non-NL-to-SQL related approaches such as direct inference using deep learning models have again been restricted to handling relational tables [66, 21, 37, 8].
Table QA benchmarks mostly focus on flat Wikipedia tables which, when compared to HCTs, are significantly simpler [64, 23, 43, 10, 9, 11, 68, 72]. FinQA  [12] consist of tables from the finance domain but specifically exclude tables with complex nested structures. There are also several benchmarks that cover only one source or domain and have a very small number of QA pairs and tables such as AIT-QA[27] with only 515 QA pairs on 116 tables and TableVQA-Bench [30] with only 250 QA pairs on 205 HCTs. Some benchmarks focus on multi-modal QA where questions are on text, tables, images or other modalities [50, 70, 69, 67, 57, 41]. Compared to these benchmarks, HCT-QA focuses on complex-layout HCTs from diverse real-world sources, provides tens of thousands of QA pairs on thousands of HCTs without requiring metadata, and provides a synthetic HCT and QA pair generator for scalability.
In Table I, we present the key distinguishing dimensions of the related benchmarks. It can be observed that each of them suffers from key limitations like a limited number of HCTs or QA pairs, a small number of questions per table, etc. Also, as the table shows, the most distinctive advantage of HCT-QA lies in its comprehensive table- and QA-level metadata, which enables fine-grained and metadata-driven analysis.
In addition, to assess the diversity of questions, we introduced two statistics: the average number of questions per table and the question syntactic similarity. The syntactic similarity is computed at the table level (for tables with more than one question) by calculating the pairwise word-level Jaccard similarity between all question pairs, followed by averaging the results. As shown, HCT-QA exhibits a higher average number of questions per table and greater lexical diversity.
Zooming-in on HiTab [13] (the closest to HCT-QA), we see that it includes HCTs from three sources and provides only 10,672 QA pairs. In contrast, HCT-QA encompasses nearly twice as many HCTs drawn from a broader range of sources and offers over seven times more QA pairs. Also, HCT-QA includes the visual representations (images) of the HCTs, thereby enabling evaluation of VLMs. Moreover, a closer examination of samples from existing benchmarks revealed several data quality issues. For instance, after manually inspecting 206 randomly selected QA samples from the HiTab benchmark, we found that 22% of the ground-truth answers were incorrect, 8% of the questions were irrelevant to their associated tables, 37% contained minor grammatical errors, and 12% included typographical mistakes. In developing HCT-QA, we placed a strong emphasis on ensuring data quality through a comprehensive and rigorous verification process.
III Real-world datasets
In this section, we present the real-world dataset. Each of its element consists of a quadruplet in the form of $(T_{HCT},Q_{NL},A,M)$, where $T_{HCT}$ is an HCT object, $Q_{NL}$ is a NL question, $A$ is the ground truth answer, and $M$ is rich metadata about $T_{HCT}$ and $Q_{NL}$. In the following, we first present the details of the HCTs collected from thousands of PDFs across diverse domains (Section III-A), and then present the details of the QA collection process (Section III-B). Figure 2 illustrates the dataset preparation process.
TABLE II: Definitions of HCT (table) properties.
Property Name Definition
Column Nesting Headers/column names are spread across more than one level (row).
  Balanced * The depth of each branch within the nesting is the same throughout the table.
  Unbalanced The depth of each branch within the nesting is not the same.
  Symmetric* Child column names repeat uniformly across table columns where nesting is present.
  Asymmetric* Child column names are different across the table in columns where nesting is present.
Row Nesting A cell spans multiple rows but is displayed as one merged cell.
  Balanced* The depth of each row branch is the same.
  Unbalanced The depth of each row branch is not the same.
  Symmetric* Child cell values repeat across the table uniformly in rows where nesting is present.
  Asymmetric* Child cell values are different across the table in rows where nesting is present.
Col. Aggregation Columns that aggregate all or a subset of other columns.
  Global* A column that aggregates all other numerical columns in the table (excluding local-group aggregation columns).
  Local* A column that aggregates a subset of numerical columns in the table.
  Explicit* An aggregation column where its name clearly states it’s an aggregation (e.g., total).
  Implicit* An aggregation column where its name is ambiguous about whether it’s an aggregation.
Row Aggregation Rows that aggregate all or a subset of other rows.
  Global* A row aggregating all other numerical rows in the table (excluding local-group aggregation rows).
  Local* A row that aggregates a subset of numerical rows in the table.
  Explicit* A row where some textual cell value clearly states it’s an aggregation (e.g., total)
  Implicit* A row where all textual cell values are ambiguous about whether it’s an aggregation.
Others
  Stand. Rel. Table* Table displays none of the other HCT properties (nesting, aggregations, labels).
III-A Real-world HCTs
@@
Data Sources. We collected thousands of PDFs from diverse sources, including Qatar National Planning Council (QNPC)  [51] Research Papers from ArXiv.org, biorxiv.org, and medarch.org, US Census covering PDF reports from the US Census Bureau [58] and Pakistan Census from the official website of the Pakistan government [48]. These sources cover various domains including education, land usage, machine learning, data management, finance, astrophysics, materials science, cryptography, astrophysics, medicine, and security. In total, we extracted 1,880 HCTs.
Extraction and Transformation. We employed the Landing AI tool [31] to generate an HTML representation of the HCTs, from which we further derived the corresponding CSV and Markdown formats. As part of the extraction pipeline, we conducted manual verification on over 10% of the extracted HCTs to ensure the tool’s ability to capture the structural complexities such as column nesting, row nesting, row groupings, and embedded aggregations. The verified sample demonstrated 100% accuracy. Also, we utilized Azure AI Document Intelligence [14] to extract HCT images. A similar 10% manual verification confirmed the absence of errors. Figure 3(a) presents statistics on each source.
Figure 2: Preparing the real-world dataset. (a) Extracting HCTs. (b) Annotating HCTs with table properties. (c) Generating QA pairs with question properties.
Characterization of HCT Properties. Across the broad range of HCTs we investigated, we observed that certain structural layouts are highly prevalent such as column nesting, row nesting, and embedded aggregations whereas others, like nested tables, occur less frequently and are considered outliers. In this study, we focus on common structural patterns and exclude HCTs exhibiting rare or irregular layouts. The dominant layout patterns (referred to as properties) are summarized in Table II. The top-level properties are highlighted in green, while additional sub-properties are specified for cases where finer distinctions are expected to influence the LLM/VLM’s comprehension of HCTs, and hence impact the QA accuracy. For instance, if an HCT exhibits column nesting (L1 in Figure 1), we further specify whether the nesting is balanced—meaning all child nodes occur at the same hierarchical depth—or unbalanced, where nesting levels vary. Referring to Figure 1, Tables (b), (c), and (d) demonstrate balanced column nesting, whereas Table (a) illustrates an unbalanced column nesting. We also record whether the nesting is symmetric—where column names and layouts are consistent across parent nodes—or asymmetric, where they differ. A similar level of detail is captured for the row nesting property (L2 in Figure 1). Embedded aggregation columns and rows also represent a common structural characteristic across HCTs. For these cases, we track whether the aggregation is global (spanning all columns or rows) or local (restricted to a subset), and whether it is explicit, e.g., “Total,” “Sum,” and “Avg.”, or implicit, e.g., “Fruits,” denoting a collection of fruit categories.
Manually creating fine-grained metadata for each HCT object represents a distinctive feature HCT-QA, enabling more in-depth analyses of LLM and VLM performance. Figure 3(a) presents aggregated statistics illustrating the distribution of tables across various property combinations. As shown, the combination of {column nesting, row aggregation, column aggregation} is the most frequent, appearing in 396 HCTs and column nesting is the most prevalent structural property overall, occurring in more than 900 HCTs. To support the systematic annotation of this metadata, we developed a custom annotation interface [18] that allows human experts to visually inspect each HCT and select the corresponding layout properties. Each annotation is subsequently reviewed and verified by a second annotator to ensure accuracy and consistency.
@@
(a) Statistics of HCTs and QA pairs (b) Main properties of real-world HCTs